# The coisotropic embedding theorem for pre-symplectic vector spaces: an alternative proof

### v2

@@ 2 A new proof of the Coisotropic Embedding Theorem
  • Construction of the symplectic structure. The symplectic structure on $\tilde{V}$ here is constructed by simply restricting $\omega^{\prime}$ to $\tilde{V}$ by taking its pull-back along $\mathfrak{i}$. Such a pull-back reads:
    $\begin{split}\mathfrak{i}^{\star}\omega^{\prime}\,=:\,\tilde{\omega}\,=\,&dx_{j}\wedge dy^{j}+\\ +&d{p_{z}}_{a}\wedge\left(\,{\rm d}z^{a}-{P_{x}}^{ja}{\rm d}x_{j}-{P_{y}}_{j}^{a}{\rm d}y^{j}\,\right)+\\ -&{p_{z}}_{a}\left(\,{\rm d}{P_{x}}^{ja}\wedge{\rm d}x_{j}+{\rm d}{P_{y}}_{j}^{a}\wedge{\rm d}y^{j}\,\right)\,,\end{split}$ (22)
    that coincides with (11). One can check that $\tilde{\omega}$ is non-degenerate only within the intersection of the image of $\mathfrak{i}$ and of the zero-section of $\mathbf{T}^{\star}V$.
  • Check of the coisotropic condition. To prove that $V$ is a coisotropic vector subspace of $(\tilde{V},\,\tilde{\omega})$, it is sufficient to compute $V^{\perp}$ and to note that it is made only by the zero vector. Consequently $V^{\perp}\subseteq V$ and, thus, $V$ is coisotropic in $(\tilde{V},\,\tilde{\omega})$.
@@ Conclusions
We provided an alternative proof of the CET for pre-symplectic vector spaces that passes through the cotangent bundle of the pre-symplectic vector space and individuates the symplectic thickening as the subspace of such a cotangent bundle selected by the Hamiltonian functions associated with the vectors in the kernel of the original pre-symplectic form. The classical ambiguity in choosing a complement for the kernel of the pre-symplectic structure is recovered as the ambiguity in embedding such a subspace into the cotangent bundle over the original pre-symplectic vector space.

### v3

@@ 2 A new proof of the Coisotropic Embedding Theorem
  • Construction of the symplectic structure. The symplectic structure is constructed by taking the pullback of the ambient symplectic form $\omega^{\prime}$ along the embedding $\mathfrak{i}$:
    $\tilde{\omega}:=\mathfrak{i}^{\star}\omega^{\prime}=\mathfrak{i}^{\star}(\omega_{\operatorname{\textup{{T}}}^{\star}M}+\rho^{\star}\omega)=\mathfrak{i}^{\star}\omega_{\operatorname{\textup{{T}}}^{\star}M}+\mathfrak{i}^{\star}\rho^{\star}\omega\,$, (21)
    which locally reads
    $\tilde{\omega}=\rho_{K^{\star}}^{\star}\omega+{\rm d}p_{a}\wedge P^{a}+p_{a}{\rm d}P^{a}\,$, (22)
    which is identical to the form (12) derived in the classical construction. Thus, as in the classical case, one can prove that $\tilde{\omega}$ is symplectic in a tubular neighborhood of the zero-section of the bundle $\tau\;:\;\;\mathbf{K}^{\star}\to M$, and that such a zero-section can be extended to the whole bundle if the connection $P$ is flat. Again, we will denote the tubular neighborhood above as $\tilde{M}$, and it represents the symplectic manifold we are searching for.
[…]
@@ Conclusions
We provided an alternative proof of the CET for pre-symplectic manifolds that passes through the cotangent bundle of the pre-symplectic manifold and identifies the symplectic thickening as the submanifold of such a cotangent bundle selected by the Hamiltonian functions associated with the vectors in the kernel of the original pre-symplectic form. The classical ambiguity in choosing a complement for the kernel of the pre-symplectic structure is recovered as the ambiguity in embedding such a submanifold into the cotangent bundle over the original pre-symplectic manifold.